### v1

# SyncLipMAE: Contrastive Masked Pretraining for Audio–Visual Talking-Face Representations

Abstract
We introduce SyncLipMAE, a self-supervised pretraining framework for talking-face video that learns synchronization-aware and transferable facial dynamics from unlabeled audio–visual streams. Our approach couples masked visual modeling with cross-modal contrastive alignment and employs three per-frame prompt tokens that explicitly encode the essential factors of a talking-face frame—identity, vocal motion (speech-synchronized facial dynamics), and ambient motion (audio-agnostic movements such as blinks and head pose). The contrastive objective uses time-aligned vocal-motion and audio tokens as positives and misaligned pairs as negatives, driving both modalities into a shared embedding space and yielding token-level audio–visual stream synchronization. After pretraining, the aligned audio tokens together with the visual prompt tokens (identity, vocal motion, ambient motion) form a unified interface for four disparate downstream settings: (i) audio–visual stream synchronization; (ii) facial emotion and head/face action recognition; (iii) visual speech recognition; and (iv) visual dubbing, for which we, for the first time, enable indistinguishable audio- or video-driven control within a single model. Across four task families that require distinct capabilities, SyncLipMAE achieves state-of-the-art results, underscoring the effectiveness of synchronization-aware, factorized self-supervised pretraining.
1 Introduction
Talking-face video underpins human–computer interaction, accessibility, content creation, and post-production. Practical systems must jointly reason about linguistic content, audio–visual timing, video editing to match a target speech track or reference motion, and facial expressions/actions. Classic synchronisation pipelines such as SyncNet and subsequent Transformer-based variants (e.g., VocaLiST) focus on detecting in-/out-of-sync pairs or removing stream lags, while recent audio–visual pretraining methods (e.g., AV-HuBERT) have shown strong transfer to lip reading and AV-ASR. Yet, these advances do not directly yield a token-level, stream-synchronised representation that cleanly separates identity, speech-synchronised mouth motion, and other facial dynamics—capabilities that are essential for both analysis and controllable generation.
Most prior work tackles one facet at a time. Visual or video masked modeling (MAE/VideoMAE) excels at reconstruction but is modality-specific and does not enforce audio–visual alignment; contrastive AV learning improves correspondence but is typically framed as binary sync classification rather than a shared token space; and dubbing systems (e.g., Wav2Lip, LatentSync, MuseTalk) optimise for generation quality and lip accuracy but are not designed to provide a unified interface that can be driven interchangeably by audio or reference motion within a single model. As a result, existing solutions often rely on wider temporal windows or task-specific heads, and they lack a factorised, alignment-aware representation that transfers seamlessly across synchronisation, understanding, and dubbing.
We introduce SyncLipMAE, a self-supervised pretraining framework tailored to talking-face video. SyncLipMAE learns three per-frame prompt tokens—identity, vocal motion (speech-synchronised dynamics), and ambient motion (audio-agnostic movements)—and couples masked visual reconstruction with a cross-modal contrastive objective that aligns vocal-motion tokens to temporally matched audio tokens in a shared embedding space. A two-view masking strategy (uniform vs. face-preserving) sources patch context and factor prompts, and a simple identity-shuffle regulariser further purifies the identity token. Decoding is performed in two passes per frame (video- and audio-conditioned) with shared weights, which provides symmetric supervision that tightens audio–visual stream synchronization at the token level. This factorised, aligned interface directly supports diverse downstream uses without task-specific architecture changes.
Beyond analysis, SyncLipMAE enables unified visual dubbing: using a WanVACE-style DiT backbone for masked video inpainting, we inject either aligned audio tokens or vocal-motion tokens after patch embedding via a lightweight AudioPack, so the same model admits no-difference audio- or video-driven control. This preserves identity and pose while re-synchronising the mouth region, and it naturally benefits from the shared token space learned during pretraining.
Contributions. (i) We propose a synchronization-aware, factorised pretraining scheme that yields three prompt tokens (identity, vocal motion, ambient motion) and a shared audio–visual token space. (ii) We introduce a two-view masking design and identity-shuffle regularisation tailored to talking faces. (iii) We show a unified dubbing interface that for the first time enables no-difference audio- or video-driven control within a single model by injecting aligned control tokens into a WanVACE backbone. (iv) We demonstrate state-of-the-art performance across four distinct tasks—audio–visual stream synchronization, facial expression/action understanding, visual speech recognition, and visual dubbing—highlighting the effectiveness and generality of the approach.
Figure 1: Panel (a) schematically illustrates the core components of SyncLipMAE and the computational pipeline used during pretraining, while panel (b) shows its adaptation to downstream tasks.
2 Related Work
Audio–visual synchronisation and correspondence.
A long line of work formulates A/V learning as correspondence or temporal alignment.  Arandjelovic & Zisserman (2017) learns generic audio–visual correspondence from unlabelled video, while  Morgado et al. (2020) casts synchronisation as in-time vs. out-of-time discrimination within the same clip. For faces,  Chung & Zisserman (2016) introduces a two-stream embedding for lip–audio alignment and lag estimation; recent transformers such as  Kadandale et al. (2022) improve robustness across speech/singing with longer-range context.
Masked modelling and contrastive learning at scale.
Contrastive pretraining aligns paired modalities at scale (e.g., CLIP  Radford et al. (2021)), while masked autoencoders (MAE He et al. (2022)/VideoMAE Tong et al. (2022)) show that reconstructing heavily masked inputs yields strong visual/video representations. These paradigms are complementary and now standard building blocks for multimodal pretraining.
Facial video representation pretraining.
Closer to our setting, MARLIN Cai et al. (2023) applies masked autoencoding to facial videos, using facial-region-guided masking to learn a universal face encoder transferable to expression recognition, deepfake detection, etc. Subsequent variants adapt MAE-style pretraining to dynamic facial expression recognition under limited labels Sun et al. (2023). While these works focus on reconstruction-only objectives within the facial domain, we couple MAE with audio–visual contrast to align speech-driven dynamics, and structure the representation to separate identity and motion factors.
Visual speech recognition and audio–visual ASR.
From early end-to-end lipreading (LipNet) to large-scale self-supervised A/V pretraining (AV-HuBERT) and automated VSR/AVSR recipe design (Auto-AVSR Ma et al. (2023)), results consistently show that robust visual streams improve recognition and benefit from pretraining on unlabelled A/V data.
Talking-face generation and lip-sync synthesis.
Audio-driven talking-head synthesis has progressed from GAN-based pipelines with explicit sync critics (e.g., Wav2Lip Prajwal et al. (2020)) to diffusion in latent/image space (e.g., LatentSync Li et al. (2024)) and efficient diffusion heads (MuseTalk Zhang et al. (2024)). In parallel, foundation video generators and unified conditional DiT Peebles & Xie (2023) frameworks (WAN Wan et al. (2025); VACE Jiang et al. (2025)) provide scalable backbones and conditioning interfaces (Video Condition Unit, context adapters) for editing or T2V, which our dubbing setup leverages.
Table 1: Lip synchronisation on Hallo3. Higher is better for Acc and R-Precision; lower is better for Offset.
Method K = 1 K = 5 K = 15 R-precision (32) Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Top1 $\uparrow$ Top2 $\uparrow$ Top3 $\uparrow$ SyncNet -5 20.37 5.04 24.18 4.03 30.03 3.02 7.47 12.77 17.72 VocaLiST -5 21.85 4.34 24.56 3.63 28.82 2.98 8.80 15.08 19.86 StableSyncNet -16 27.95 2.78 28.60 2.72 31.66 2.72 14.40 25.65 34.52 SyncLipMAE  -1 (ours) 52.53 2.66 68.41 1.73 82.27 0.93 40.18 61.48 73.49
3 SyncLipMAE: Pretraining Objective and Architecture
3.1 Overview and Design Goals
SyncLipMAE aims to learn synchronization-aware talking-face representations from unlabeled audio–visual streams. As illustrated in figure 1(a), our approach couples MAE-style masked image/video modeling on the visual stream with a CLIP-style cross-modal contrastive alignment that treats time-aligned vocal-motion and audio tokens as positives and temporally misaligned pairs as negatives, driving both modalities into a shared embedding space and yielding token-level audio–visual stream synchronization (§3.3). Per frame, a ViT-based encoder exposes three prompt tokens that explicitly encode the essential factors of a talking-face frame—identity $\mathbf{z}^{\mathrm{id}}$, vocal motion $\mathbf{z}^{\mathrm{voc}}$, and ambient motion $\mathbf{z}^{\mathrm{amb}}$—and the decoder reconstructs frames via cross-attention conditioned on these prompts.
Inputs and Notation.
Let $\mathbf{x}_{1:T}\in\mathbb{R}^{T\times C\times H\times W}$ be video frames and $\mathbf{a}\in\mathbb{R}^{N}$ mono audio at sampling rate $f_{s}$. Each frame is patchified into $N_{p}$ patches and embedded to $D$-dimensional tokens. For MAE-style masking, denote by $\mathcal{V}_{t}\subset\{1,\dots,N_{p}\}$ the indices of visible patches at time $t$ and by $\mathcal{M}_{t}$ the masked ones, with $|\mathcal{V}_{t}|=N_{\mathrm{vis}}$ and $\mathcal{V}_{t}\cup\mathcal{M}_{t}=\{1,\dots,N_{p}\}$. The per-frame visible patch tokens are $\mathbf{P}^{\mathrm{vis}}_{t}\in\mathbb{R}^{N_{\mathrm{vis}}\times D}$. Let $\mathbf{e}^{\mathrm{mask}}\in\mathbb{R}^{D}$ be a learnable mask token; a restoration operator places $\mathbf{P}^{\mathrm{vis}}_{t}$ and copies of $\mathbf{e}^{\mathrm{mask}}$ at indices $\mathcal{M}_{t}$ to form a length-$N_{p}$ sequence for decoding. Per frame, the encoder also outputs three prompt tokens, $\mathbf{z}^{\mathrm{id}}_{t},\ \mathbf{z}^{\mathrm{amb}}_{t},\ \mathbf{z}^{\mathrm{voc}}_{t}\in\mathbb{R}^{D}$. For audio, we employ a pretrained speech encoder with $L$ layers that is kept frozen during pretraining; let $\{\mathbf{H}^{(\ell)}\}_{\ell=1}^{L}$ denote the hidden-state sequences from all layers. To obtain frame-wise alignment with the video, we resample the input waveform so that the encoder emits $T$ tokens per utterance (one token per video frame) at every layer, yielding $\mathbf{A}^{(\ell)}=\{\mathbf{a}^{(\ell)}_{t}\}_{t=1}^{T}$.
3.2 Model Design
Two-Bypass Face-Aware Masking.
[…]
Table 2: Lip synchronisation on VFHQ.
Method K = 1 K = 5 K = 15 R-precision (32) Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Top1 $\uparrow$ Top2 $\uparrow$ Top3 $\uparrow$ SyncNet -5 22.16 5.14 26.47 4.34 31.09 3.29 9.40 16.91 23.09 VocaLiST -5 21.76 5.14 24.34 4.57 29.23 3.62 10.18 16.56 21.98 StableSyncNet -16 32.40 3.17 33.06 3.07 35.43 2.90 23.54 37.65 46.77 SyncLipMAE -1 (ours) 38.58 3.80 48.95 3.06 61.78 2.16 35.52 52.98 63.37
Visual Encoder and Prompt Tokens.
A ViT-style encoder applied to the visible patches outputs per-frame patch tokens (from Bypass 1) and three prompt tokens
$\mathbf{Z}^{\text{prompt}}_{t}=\big[\mathbf{z}^{\text{id}}_{t},\ \mathbf{z}^{\text{amb}}_{t},\ \mathbf{z}^{\text{voc}}_{t}\big]\in\mathbb{R}^{3\times D},$
where $\mathbf{z}^{\mathrm{id}}_{t}$ is taken from Bypass 1 and $\mathbf{z}^{\mathrm{voc}}_{t},\mathbf{z}^{\mathrm{amb}}_{t}$ from Bypass 2 as defined above.
Identity-Consistent Prompt Shuffling.
Within a mini-batch, to promote motion-invariant identity embeddings, we randomly swap $\mathbf{z}^{\mathrm{id}}$ among frames that share the same subject identity but exhibit different facial motions, weakening incidental coupling between identity prompts and short-term dynamics.
Audio Features and Adapter.
A pretrained $L$-layer speech encoder (kept frozen) produces per-layer hidden sequences; we resample the input waveform so that each layer emits $T$ tokens (one per video frame), yielding aligned streams $\{\mathbf{a}^{(\ell)}_{t}\}_{t=1}^{T}$. To preserve both low-level acoustic detail and higher-level linguistic cues, we concatenate the aligned per-layer tokens along the feature dimension at each time step,
$\tilde{\mathbf{a}}_{t}=\big[\mathbf{a}^{(1)}_{t}\,\|\,\cdots\,\|\,\mathbf{a}^{(L)}_{t}\big$],
and pass them through an audio adapter that projects to the shared width $D$, producing $\mathbf{A}_{t}\in\mathbb{R}^{D}$. This design mitigates the bias of top encoder layers toward semantic content induced by ASR pretraining while retaining waveform-proximal information from earlier layers.
Decoder with Prompt Cross-Attention.
We use an MAE-style decoder with $N_{\text{dec}}$ Transformer blocks. Its input at each frame is the restored token sequence formed by encoded visible patches from Bypass 1 plus mask tokens (with positional embeddings). Each block first applies self-attention over this full sequence (visible patches + mask tokens), and then cross-attends to the prompt tokens—identity, ambient, and a conditioning token $\mathbf{c}_{t}$. We perform two decoding passes per frame that share decoder weights: (i) a video-driven pass with $\mathbf{c}_{t}=\mathbf{z}^{\mathrm{voc}}_{t}$ producing $\hat{\mathbf{x}}^{\mathrm{voc}}_{t}$, and (ii) an audio-driven pass with $\mathbf{c}_{t}=\mathbf{A}_{t}$ producing $\hat{\mathbf{x}}^{\mathrm{aud}}_{t}$. Both passes reconstruct the same target frame $\mathbf{x}_{t}$, providing symmetric supervision that encourages the vocal-motion and audio tokens to encode consistent, alignable information. A linear head maps decoded tokens back to pixels.
3.3 Pretraining Objectives
Pixel-Space Reconstruction (MSE).
We minimize per-frame mean squared error for the two decoding passes (video-driven and audio-driven; see §3.2):
$\mathcal{L}_{\text{pix}}^{\text{voc}}=\tfrac{1}{T}\sum_{t=1}^{T}\!\big\|\hat{\mathbf{x}}^{\text{voc}}_{t}-\mathbf{x}_{t}\big\|_{2}^{2},\qquad\mathcal{L}_{\text{pix}}^{\text{aud}}=\tfrac{1}{T}\sum_{t=1}^{T}\!\big\|\hat{\mathbf{x}}^{\text{aud}}_{t}-\mathbf{x}_{t}\big\|_{2}^{2}.$
Audio–Vocal Contrastive Alignment.
Unlike prior audio–visual synchronisation systems that cast the task as in-/out-of-sync binary classification or as pairwise contrastive matching, we align per-frame audio tokens $\mathbf{A}_{t}$ with vocal-motion tokens $\mathbf{v}_{s}=\mathbf{z}^{\mathrm{voc}}_{s}$ using a CLIP-style symmetric InfoNCE. Let $p_{t\to s}=\frac{\exp(\langle\mathbf{A}_{t},\mathbf{v}_{s}\rangle/\tau)}{\sum_{u=1}^{T}\exp(\langle\mathbf{A}_{t},\mathbf{v}_{u}\rangle/\tau)}$ be the softmax over vocal candidates for audio index $t$, and $P_{t}$ the positive set (aligned index and optional neighbors within $\pm k$, expanded by an audio self-similarity threshold). The loss is
$\mathcal{L}_{\mathrm{CL}}=-\frac{1}{2T}\sum_{t=1}^{T}\Big[\log\textstyle\sum_{s\in P_{t}}p_{t\to s}\;+\;\log\textstyle\sum_{s\in P_{t}}p_{s\to t}\Big],$
with temperature $\tau$. This preserves CLIP-style bidirectional alignment while allowing multiple positives per anchor.
Cross-Covariance Decorrelation.
To reduce leakage across factor-specific embeddings, we penalize cross-covariance between centered tokens. For $\mathcal{P}=\{\mathrm{id},\,\mathrm{amb},\,\mathrm{voc}\}$:
$\mathcal{L}_{\text{cov}}=\frac{1}{D^{2}}\sum_{\{p,q\}\subset\mathcal{P}}\big\|\mathrm{Cov}\!\big(\mathbf{z}^{p},\mathbf{z}^{q}\big)\big\|_{F}^{2},$
where covariance is computed after centering along the batch/time axis.
Total Objective.
The overall loss is
$\mathcal{L}=\lambda_{\text{pix}}\big(\mathcal{L}_{\text{pix}}^{\text{voc}}+\mathcal{L}_{\text{pix}}^{\text{aud}}\big)+\lambda_{\text{CL}}\,\mathcal{L}_{\text{CL}}+\lambda_{\text{cov}}\,\mathcal{L}_{\text{cov}}.$
Table 3: CelebV-HQ emotion and action classification. Higher is better (Accuracy/AUC).
Method Frames Emotion Action Accuracy $\uparrow$ AUC $\uparrow$ Accuracy $\uparrow$ AUC $\uparrow$ VideoMAE $1\times 16$ 68.63 60.32 93.97 84.31 VideoMAE $5\times 16$ 68.63 66.60 94.54 86.01 MARLIN $1\times 16$ 66.67 58.15 94.69 82.26 MARLIN $5\times 16$ 69.61 64.98 94.86 83.90 SyncLipMAE (ours) 16 76.47 73.94 95.06 86.45 SyncLipMAE (ours) 80 77.45 78.33 95.11 86.69
4 Downstream Adaptation
Table 4: Video dubbing on VFHQ and Hallo3. A- denotes audio-driven control; V- denotes video-driven control. Higher is better for $\mathrm{Sync}_{\mathrm{conf}}$; lower is better for FID/FVD.
Method VFHQ Hallo3 $\mathrm{Sync}_{\mathrm{conf}}\,\uparrow$ FID $\downarrow$ FVD $\downarrow$ $\mathrm{Sync}_{\mathrm{conf}}\,\uparrow$ FID $\downarrow$ FVD $\downarrow$ MuseTalk Zhang et al. (2024) 0.6820 24.44 175.16 0.6904 13.80 86.46 LatentSync-1.5 Li et al. (2024) 0.7681 17.92 131.04 0.7725 10.59 78.89 A-SyncLipMAE (ours) 0.7830 16.46 100.44 0.8027 10.04 74.01 V-SyncLipMAE (ours) 17.22 109.91 10.82 77.57
We apply SyncLipMAE to four tasks: (i) audio–visual stream synchronization — infer frame-level lag and synchrony by comparing aligned audio tokens $\mathbf{A}_{t}$ with vocal-motion tokens $\mathbf{z}^{\mathrm{voc}}_{t}$; (ii) facial understanding — infer emotion and head/face actions from the vocal- and ambient-motion tokens; (iii) visual speech recognition (VSR) — infer text from the vocal-motion tokens; (iv) visual dubbing — edit a source face video to match target speech or a reference motion, supporting audio- or video-driven control within a single model.
4.1 Audio–visual synchronisation
We estimate audio–video lag directly in the pretrained shared space without any extra head or training. Given per-frame audio tokens $\mathbf{A}_{t}$ and vocal-motion tokens $\mathbf{z}^{\mathrm{voc}}_{s}$, compute frame-wise cosine similarity
$s(t,s)=\langle\mathbf{A}_{t},\mathbf{z}^{\mathrm{voc}}_{s}\rangle,$
and aggregate along temporal offsets
$S(\Delta)=\tfrac{1}{T-|\Delta|}\sum_{t=1}^{T-|\Delta|}s\!\big(t,\,t{+}\Delta\big),\qquad\hat{\Delta}=\arg\max_{\Delta\in[-\Delta_{\max},\,\Delta_{\max}]}S(\Delta).$
The in-sync score is $S(0)$ and the estimated lag is $\hat{\Delta}$. All computations use the frozen tokens from pretraining; no correspondence MLP or additional supervision is introduced.
4.2 Facial understanding: emotion and action
We use SyncLipMAE ’s motion factors per frame and keep the prediction head minimal. For each frame $t$, we form a motion descriptor by averaging the two factors, $\mathbf{u}_{t}=\tfrac{1}{2}\big(\mathbf{z}^{\mathrm{voc}}_{t}+\mathbf{z}^{\mathrm{amb}}_{t}\big)$, and apply a single linear classifier to obtain frame-level logits. Video-level logits are the mean of frame-level logits over the sampled frames; the final prediction is the softmax over the aggregated logits. This setup probes whether the representation cleanly separates speech-synchronized orofacial motion from audio-agnostic dynamics (e.g., blinks, head pose) while keeping the adaptation lightweight; datasets such as CelebV-HQ provide the emotion/action labels used in this setting.
4.3 Lip reading (VSR)
We use a visual-only sequence-to-sequence head on top of SyncLipMAE ’s vocal-motion tokens. Let $\mathbf{V}=\{\mathbf{z}^{\mathrm{voc}}_{t}\}_{t=1}^{T^{\prime}}$ be the input to a Conformer encoder, yielding hidden states $\mathbf{H}=\{\mathbf{h}_{t}\}_{t=1}^{T^{\prime}}$. A Transformer decoder autoregressively predicts subword units with cross-attention to $\mathbf{H}$.
Losses. Let the target token sequence be $Y=\{y_{u}\}_{u=1}^{U}$ over vocabulary $\Sigma$ (blank $\varnothing$ for CTC). The CTC branch defines
$P_{\mathrm{CTC}}(Y\,|\,\mathbf{H})=\!\!\sum_{\pi\in\mathcal{B}^{-1}(Y)}\ \prod_{t=1}^{T^{\prime}}p_{\mathrm{ctc}}(\pi_{t}\,|\,\mathbf{h}_{t}),\qquad\mathcal{L}_{\mathrm{CTC}}=-\log P_{\mathrm{CTC}}(Y\,|\,\mathbf{H}),$
where $\mathcal{B}$ collapses repeats and removes blanks. The decoder branch is trained with teacher forcing:
$\mathcal{L}_{\mathrm{DEC}}=-\sum_{u=1}^{U}\log p_{\mathrm{dec}}\!\left(y_{u}\,\middle|\,y_{<u},\,\mathbf{H}\right).$
The total objective is the standard hybrid form
$\mathcal{L}_{\mathrm{VSR}}=\lambda_{\mathrm{CTC}}\ \mathcal{L}_{\mathrm{CTC}}+(1-\lambda_{\mathrm{CTC}})\ \mathcal{L}_{\mathrm{DEC}},$
with a tunable weight $\lambda_{\mathrm{CTC}}\in(0,1)$.
4.4 Video dubbing
We cast dubbing as masked video inpainting on face crops within a VACE-style unified interface: task inputs (reference frames and a spatiotemporal mouth mask) are organized as a Video Condition Unit and processed by a Diffusion-Transformer backbone with Concept Decoupling and Context Adapters, following VACE. Concretely, we instantiate the pipeline on WanVACE and keep its architecture, losses, and training schedule unchanged. WanVACE natively accepts multiple reference images (1–3 in the public release), exposed as a comma-separated list, which we leverage for identity control.
Control injection. Inspired by audio-conditioned injection in recent avatar generation, we adopt a minimal AudioPack placed after the VACE patch-embedding stage and before the first DiT VACE block. Per-frame control tokens—either audio tokens $\mathbf{A}_{t}$ produced by SyncLipMAE or reference vocal-motion tokens $\mathbf{z}^{\mathrm{voc}}_{t}$—are temporally encoded and projected to the DiT width, then added to the patch-embedded video token stream. No other WanVACE components are modified; switching $\mathbf{A}_{t}$ vs. $\mathbf{z}^{\mathrm{voc}}_{t}$ toggles audio-driven vs. video-driven control within the same model.
Long-form continuity via dual references. To support arbitrarily long, coherent dubbing, we provide two references: (i) an identity image (stable appearance prior), and (ii) the last frame of the previous segment (temporal bridge). During training we randomly drop the continuity reference with 10% probability, so the model can also initialize the first segment without a history frame.
Masking strategy. Following common lip-sync inpainting practice, we detect the facial region (chin to hairline) and mask the lower half as the editable area; to accommodate jaw excursions (post-edit lip motion often increases mouth opening), the mask is expanded slightly downward beyond the chin. This aligns with widely used lower-face inpainting setups in Wav2Lip-style and diffusion-based lip-sync systems.
[…]
Figure 2: Qualitative visual comparison: LatentSync vs. our approach.
5 Experiments
Implementation details.
Backbones. The visual encoder follows the Sapiens-0.3B configuration and is initialized from its pretrained weights; the decoder adopts the ViTMAE-B design with a cross-attention sublayer in every block to attend to the identity/ambient/vocal prompt tokens. The audio branch uses a pretrained wav2vec 2.0 base encoder kept frozen during pretraining. Pretraining. SyncLipMAE is trained on 128 GPUs with a per-GPU batch size of 48 (global $128{\times}48$) for 400k steps ( 15 days). Video dubbing fine-tuning. We initialize from WanVACE-1.3B, freeze Wan backbone blocks, and optimize only VACE blocks ( 0.7B trainable params). Training uses 128 GPUs, batch size 1 per GPU, for  30k steps ( 3 days). Each sample contains 17+1+117{+}1{+}1 frames: 17 target frames, one history frame (for temporal continuity), and one identity frame (appearance prior).
Pretraining data.
We pretrain SyncLipMAE on a mixture of talking-face corpora: Hallo3 Cui et al. (2025), CelebV-HQ Zhu et al. (2022), CelebV-Text Yu et al. (2023), MEAD Wang et al. (2020), VFHQ Xie et al. (2022), HDTF Zhang et al. (2021) and RAVDESS Livingstone & Russo (2018). Unless specified, audio is mono 16 kHz and videos are 25 fps. We adopt the official train–val–test split when available; otherwise, we randomly sample 100 instances each for validation and testing, and use the remainder for training.
Audio–visual stream synchronization.
[…]
Temporal lag detection. For a candidate lag $\tau\!\in\!\mathcal{T}$ (positive: video lags audio), define the average distance
$D(\tau)=\frac{1}{T_{\tau}}\sum_{t=1}^{T_{\tau}}d\!\big(a_{t},\,v_{t+\tau}\big),$
with $T_{\tau}$ chosen so indices are valid and $d(\cdot,\cdot)$ a feature distance (e.g., cosine or $\ell_{2}$). The estimated lag is
$\hat{\tau}=\arg\min_{\tau\in\mathcal{T}}D(\tau).$
Given ground-truth lag $\tau^{\star}$, we report
$\mathrm{Offset}=\tfrac{1}{N}\sum_{s=1}^{N}|\hat{\tau}^{(s)}-\tau^{(s)}|\quad\text{and}\quad\mathrm{Acc}_{\pm K}=\tfrac{1}{N}\sum_{s=1}^{N}\mathbf{1}\!\big(|\hat{\tau}^{(s)}-\tau^{(s)}|\leq K\big).$
[…]
$\mathrm{R\text{-}precision@}k\;=\;\frac{1}{B}\sum_{i=1}^{B}\mathbf{1}\big(\mathrm{rank}_{i}(v_{i})\leq k\big),$
reported for small $k$ (e.g., k=1,2,3k\!=\!1,2,3) and fixed pool size (e.g., 32).
Results.  Tab 1 and  Tab 2 summarise results on Hallo3 and VFHQ. Across both datasets, SyncLipMAE consistently improves temporal alignment (higher $\mathrm{Acc}_{\pm K}$, lower Offset) and cross-modal retrieval (higher R-precision@k). Notably, these gains are achieved with single-frame visual conditioning (n=1n{=}1), whereas baselines rely on wider temporal windows (n=5n{=}5 or $16$), indicating stronger per-frame audio–visual correspondence and reduced dependence on temporal aggregation.
Facial understanding (emotion & actions).
We evaluate on CelebV-HQ Zhu et al. (2022) using the emotion and action attribute sets, reporting video-level Accuracy and AUC (higher is better). Emotion is treated as a single-label, 8-way classification task, whereas action is a multi-label problem over 35 categories (one-vs-rest scoring aggregated at the video level). To match common practice, clip-based baselines (e.g., VideoMAE/MARLIN) adopt the $a{\times}b$ protocol—uniformly sample $a$ clips of $b$ frames and average their predictions following  Tong et al. (2022); frame-averaging models use $n$ uniformly sampled frames and average per-frame predictions Korbar et al. (2019). All methods use a single linear classifier on frozen features. Let $\{\mathbf{x}_{i}\}_{i=1}^{m}$ be the sampled clips (or frames), $f(\cdot)$ the frozen encoder, and $W$ the linear head; the video-level logits are
$\bar{\ell}(V)=\frac{1}{m}\sum_{i=1}^{m}W\,f(\mathbf{x}_{i}),\qquad\hat{y}(V)=\arg\max_{k}\mathrm{softmax}_{k}\!\big(\bar{\ell}(V)\big),$
with Accuracy/AUC computed from $\hat{y}(V)$ at the video level.
Results. Tab 3 summarises CelebV-HQ emotion and action. On emotion, SyncLipMAE clearly surpasses the best baseline with 16 frames and improves further at 80 frames, indicating benefits from longer temporal context. On action—where baselines are already strong—SyncLipMAE provides consistent but modest gains, with a slight boost when using 80 frames. Overall, SyncLipMAE delivers substantial improvements on emotion and steady gains on action.
Visual Speech Recognition (VSR)
We evaluate SyncLipMAE on HDTF, VFHQ, and Hallo3, predicting text from silent videos. Following standard practice, accuracy is measured by Word Error Rate (WER), computed from word-level substitutions ($S$), deletions ($D$), and insertions ($I$) against the reference of length $N$:
$\mathrm{WER}=\frac{S+D+I}{N}.$
Results. Tab 5 reports WER on VFHQ, Hallo3, and HDTF. The visual-only head built on SyncLipMAE ’s tokens achieves the best WER on all three sets: $7.37$ on VFHQ, $5.99$ on Hallo3, and $5.78$ on HDTF, improving over Auto-AVSR by absolute $0.76$, $0.24$, and $0.13$ respectively, and substantially outperforming AV-HuBERT on each set. These gains indicate that SyncLipMAE ’s pretraining yields visual motion features that transfer effectively to sentence-level lip reading under the standard hybrid CTC/attention recipe.
Video dubbing.
We evaluate video dubbing—re-synchronising the mouth region of a source video to a target speech while preserving identity, pose, and expressions—using a WanVACE-based model that supports both audio- and video-driven conditioning via a lightweight vocal-cue injection (details in section 4.4). We report the metrics in Tab 4. $\mathrm{Sync}_{\mathrm{conf}}$ (identical to LSE-C) is computed with the stable SyncNet evaluator from LatentSync Li et al. (2024), replacing the commonly used vanilla SyncNet whose scores are less reliable in our experiments Prajwal et al. (2020). Following SyncNet Chung & Zisserman (2016), for each short sliding window we form an audio–video distance curve over temporal offsets $\Delta\!\in\![-L,L]$:
$d_{t}(\Delta)\;=\;\big\|\,\phi_{a}\!\big(a_{t:t+w-1}\big)\;-\;\phi_{v}\!\big(v_{t+\Delta:\,t+\Delta+w-1}\big)\,\big\|_{2},$
and define the per-window confidence as the median–minimum gap on this curve,
$\mathrm{Conf}_{t}\;=\;\mathrm{median}_{\Delta}\,d_{t}(\Delta)\;-\;\min_{\Delta}\,d_{t}(\Delta),\qquad\mathrm{Sync}_{\mathrm{conf}}\;=\;\frac{1}{T}\sum_{t=1}^{T}\mathrm{Conf}_{t},$
where larger values indicate a clearer minimum at the correct offset (stronger A/V synchrony) Chung & Zisserman (2016); Prajwal et al. (2020). FID measures per-frame perceptual quality via the Fréchet distance between Inception features of real vs. generated frames (lower is better) Heusel et al. (2017); FVD extends this to spatiotemporal video features to capture visual quality and temporal coherence jointly (lower is better) Unterthiner et al. (2018).
Results.
Qualitatively (Fig. 2), SyncLipMAE preserves identity and fine lip details (e.g., teeth, beard) and is more robust to occlusions than LatentSync/MuseTalk. Quantitatively (Tab. 4), consistent with these observations, the audio-driven variant (A-SyncLipMAE) attains the best overall scores on synchrony ($\mathrm{Sync}_{\mathrm{conf}}$) and perceptual quality (FID/FVD) on both VFHQ and Hallo3. The video-driven variant (V-SyncLipMAE) closely matches the audio-driven performance and surpasses prior methods, indicating that the shared token space indeed aligns modalities and enables no-difference audio- or video-driven control within a single model.
Ablation Studies.
We perform ablations on the key components of SyncLipMAE, examining: (i) whether to enable Two-Bypass Face-Aware Masking; (ii) the strategy for adapting audio features; and (iii) the inclusion of cross-attention in the decoder. Detailed results are provided in the Appendix.
6 Conclusion
We introduced SyncLipMAE, a self-supervised framework for talking-face video that couples masked visual reconstruction with contrastive audio–visual alignment and explicitly factorizes each frame into three prompt tokens—identity, vocal motion, and ambient motion. This design yields synchronization-aware, disentangled features that transfer directly to diverse applications. Across four tasks—audio–visual stream synchronization, facial understanding, visual speech recognition, and unified visual dubbing—SyncLipMAE achieves state-of-the-art performance, validating the effectiveness of this pretraining strategy.
Appendix A Ablation Studies
Table 6: Ablations on Hallo3 (audio–visual stream synchronization). Higher is better for Acc and R-Precision; lower is better for Offset.
Variant K = 1 K = 5 K = 15 R-precision (32) Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Top1 $\uparrow$ Top2 $\uparrow$ Top3 $\uparrow$ (A) Two-Bypass Face-Aware Masking A1  Uniform-only (single view) 9.34 7.18 12.68 6.22 18.57 5.36 2.83 5.14 7.92 A2  Face-aware-only (single view) 45.63 3.45 61.07 2.37 75.88 1.27 33.45 52.81 66.19 A3  Two-bypass w/o photometric 49.62 3.02 65.91 2.03 79.87 1.05 37.54 57.89 70.86 (B) Audio Feature Adaptation B1  Last-layer only 42.68 3.62 58.34 2.51 73.45 1.36 30.92 49.87 63.11 (C) Decoder Cross-Attention C1  No cross-attention 48.27 3.10 64.31 2.09 78.44 1.09 36.21 56.48 69.32 Ours 52.53 2.66 68.41 1.73 82.27 0.93 40.18 61.48 73.49
Table 7: Audio–visual stream synchronization on CelebV-HQ. Higher is better for Acc/R-precision; lower is better for Offset.
Smooth window K=1 K=5 K=15 R-precision (32) Method Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Top1 $\uparrow$ Top2 $\uparrow$ Top3 $\uparrow$ SyncNet 20.26 6.14 24.50 5.67 32.07 4.99 8.08 14.90 20.91 VocaLiST – LRS2 17.42 6.60 20.46 6.38 25.96 5.89 5.98 11.59 16.45 StableSyncNet 32.95 4.66 34.35 4.62 38.04 4.44 21.26 33.20 40.69 Ours 32.63 4.83 40.96 4.41 52.91 3.56 28.47 44.19 54.32
Table 8: Audio–visual stream synchronization on CelebV-Text. Higher is better for Acc/R-precision; lower is better for Offset.
Smooth window K=1 K=5 K=15 R-precision (32) Method Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Top1 $\uparrow$ Top2 $\uparrow$ Top3 $\uparrow$ SyncNet 18.80 6.10 22.73 5.40 28.85 4.56 8.89 15.03 20.54 VocaLiST – LRS2 22.05 5.87 25.96 5.39 32.27 4.64 11.08 17.66 23.04 StableSyncNet 34.42 3.93 35.51 3.88 37.58 3.76 27.02 37.25 43.35 Ours 34.95 5.10 46.46 4.44 60.60 3.51 25.80 39.02 47.99
Table 9: Audio–visual stream synchronization on MEAD. Higher is better for Acc/R-precision; lower is better for Offset.
Smooth window K=1 K=5 K=15 R-precision (32) Method Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Top1 $\uparrow$ Top2 $\uparrow$ Top3 $\uparrow$ SyncNet 39.39 4.01 50.03 2.97 61.18 1.85 20.74 32.29 40.01 VocaLiST – LRS2 35.02 4.59 43.81 3.64 55.60 2.50 15.69 26.03 33.61 StableSyncNet 57.51 2.05 60.36 1.85 64.25 1.57 37.20 54.19 62.83 Ours 45.90 2.77 56.20 2.04 69.76 1.30 38.63 60.94 74.36
Table 10: Audio–visual stream synchronization on RAVDESS. Higher is better for Acc/R-precision; lower is better for Offset.
Smooth window K=1 K=5 K=15 R-precision (32) Method Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Top1 $\uparrow$ Top2 $\uparrow$ Top3 $\uparrow$ SyncNet 20.54 9.38 26.76 8.79 38.52 7.54 14.47 22.69 28.48 VocaLiST – LRS2 22.12 8.93 27.97 8.33 39.96 7.15 14.94 23.46 29.21 StableSyncNet 42.69 6.91 46.06 6.62 53.79 5.97 34.51 46.21 51.28 Ours 47.94 3.09 61.03 2.37 79.03 1.23 39.97 62.90 76.26
Two-Bypass Face-Aware Masking.
We evaluate four variants on Hallo3 using the same A/V synchronisation protocol as the main results (Acc±K, Offset, R-precision; see Tab 6). A1 (Uniform-only) severely underperforms: heavy uniform masking hides the lip region, the contrastive loss scarcely decreases, and all sync metrics degrade markedly. A2 (Face-aware-only) trains the contrastive objective but yields weaker reconstruction guidance, producing normal yet inferior sync scores. A3 (Two-bypass w/o photometric) restores learning stability but, without motion-view photometric jitter, motion tokens leak appearance; alignment improves over A2 but remains below the best. A4 (Two-bypass + photometric, ours) is strongest across all metrics, confirming that a uniform view for reconstruction context plus a face-preserving view—with photometric perturbations—for prompt tokens best supports token-level A/V alignment.
Audio Feature Adaptation.
[…]
Decoder Cross-Attention.
[…]
Appendix B More Results
Audio–visual stream synchronization.
We additionally evaluate SyncLipMAE on CelebV-HQ, CelebV-Text, RAVDESS, and MEAD, and report these results in the Appendix. We do so because these corpora contain substantial non-speech or highly repeated content that can confound lag metrics and retrieval: CelebV-HQ and CelebV-Text are broad, face-centric video sets with attributes beyond active speech (appearance, actions, emotions, diverse in-the-wild clips), so many segments are not strictly speaking-focused; RAVDESS and MEAD are acted emotion datasets captured in controlled settings, with RAVDESS using only two fixed sentences across many takes (plus sung versions), leading to heavy lexical repetition.
Experimental analysis. Across all four datasets, SyncLipMAE shows consistent gains in both temporal alignment (higher $\mathrm{Acc}_{\pm K}$, lower Offset) and cross-modal retrieval (higher R-precision@k). Improvements are most pronounced on the controlled emotion corpora (MEAD/RAVDESS), where reduced acoustic and visual variability amplifies the benefit of our token-level alignment; on the in-the-wild CelebV-HQ/CelebV-Text, gains remain steady despite non-speech spans and repeated content. Notably, all results are achieved with single-frame visual conditioning (n=1n{=}1), indicating strong per-frame audio–visual correspondence without reliance on long temporal windows.
Appendix C Limitation
While SyncLipMAE achieves strong results across four disparate tasks, several limitations remain. (1) Model size and deployability. Compared with classic synchronisation backbones such as SyncNet—which is lightweight enough to be used directly as an evaluation head or even as a discriminator/loss in generation systems like Wav2Lip—SyncLipMAE is substantially heavier, making “drop-in loss” usage during training less practical and raising compute and memory costs for deployment.
(2) Scope of applicability. Although SyncLipMAE aligns audio and visual tokens well for lip-sync, expression/action understanding, and VSR, it is not a universal solution for all audio–face problems. For example, when we followed a speech-to-3DMM pipeline (in the spirit of audio-driven 3D talking-head methods) and drove 3D parameters using SyncLipMAE ’s audio branch, the results were generally underwhelming—suggesting that additional geometry-aware supervision or models tailored to 3D priors are needed. (3) Factorization remains preliminary. Our decomposition into identity, vocal motion, and ambient motion is a first step; most of our downstream usage centers on the vocal-motion token. Systematically exploiting the other two factors (e.g., identity-aware editing, ambient-motion analysis/transfer) and strengthening disentanglement are promising directions for future work.